# A 4 - 8 GHz Kinetic Inductance Travelling-Wave Parametric Amplifier Using Four-Wave Mixing with Near Quantum-Limit Noise Performance

### v3

@@ IV Conclusion
In conclusion, we have designed, fabricated, and measured a wide-bandwidth Four-Wave Mixing KI-TWPA that produces gain between 3 - 9 GHz with a 1 dB compression of -58 dBm suitable for the readout of large detector arrays and superconducting qubits. Using a Y-factor method, we have demonstrated a near quantum-limited noise performance of KI-TWPA in the 4.6 to 8.1 GHz frequency range. The amplifier produces gain in disjointed frequency ranges that we call the signal band and a higher-frequency idler (or image) band. This makes it straightforward to separate the idler using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies for the readout of cryogenic detectors. The signal band peak gain of this device is shown to be tunable with the frequency of the pump tone. This tunability makes the KI-TWPA suitable for applications where the signal frequency range needs to be adjusted, such as in dark matter experiments. In addition, The higher frequency idler band of the KI-TWPA can be used to look for dark matter candidates with a higher mass range, which makes it possible to cover a wide range of masses with a single KI-TWPA.
@@ Acknowledgements.
Figure 5: a) Circuit diagram of the noise measurement setup. The input line is attenuated using three cryogenic attenuators and then connected to switch C1. Two cryogenic 50 $\mathrm{\Omega}$ loads were used as noise sources. The ‘Hot’ noise source was placed at the 3 K stage of the Dilution fridge, and the ‘Cold’ noise source was placed at the MXC. In this configuration, we can switch between the 50 $\mathrm{\Omega}$ for Y-factor measurements and the signal line for gain measurements. The Pump tone was injected in the KI-TWPA through a bandpass filter and DPX 1 and 2 with cross-over frequencies of 14 and 9 GHz, respectively. The pump tone is then routed out using DPX 3 (cross-over frequency of 9 GHz) and terminated on the 50 mK stage of the cryostat to prevent the HEMT from saturating. The cables used in the measurement setup connecting switch C1 to the ‘hot’ and ‘cold’ terminations on the 3 K and the MXC stages were NbTi superconducting cables. To calibrate the loss in the components, we used a cryogenic cable between switches B1 and B2 to bypass the TWPA and the diplexers. The output is then read out through an isolator, a cryogenic HEMT amplifier, and a room-temperature amplifier using a spectrum analyzer. b) The plot of measured transmission through the entire noise setup with the KI-TWPA unpumped is denoted as ‘Pump off,’ and through the cryogenic cable bypassing the components is denoted as ‘Bypass.’ The transmission plot labeled ‘Loss’ is the difference between the two measurements divided by two to account for the loss between switch B1 and the KI-TWPA.
@@ Appendix A Noise Measurement Setup
The pump input was filtered using a bandpass filter at the MXC stage and directed to the KI-TWPA using DPX 2 and 1 as illustrated in Fig 5a. Using the highpass port of DPX 3, the pump tone is terminated at the 50 mK stage of the dilution fridge. The total transmission of the system through the device was measured using a VNA with the pump tone off and the transmission of the cable bypassing the KI-TWPA and the diplexers as shown in Fig. 5.a. The plots comparing the transmission between the cryogenic cable and the unpumped KI-TWPA are shown in Fig. 5.b and labeled as ‘Bypass’ and ‘Pump off,’ respectively. The net loss from DPX 1 placed between the KI-TWPA and switch C1 in the circuit is also plotted in Fig. 5.b as ‘Loss.’
@@ Appendix B Noise Theory
To model the noise in the setup, we used the circuit diagram shown in Fig. 4.b. If we assume the mK losses in the components between the noise sources (hot & cold in this case) and the KI-TWPA are $L_{1}$ and $L_{2}$, the cascaded noise equations can be written as follows:
@@
$N_{tot}$ $=G_{w}N_{4}\ .$ (10)
Where $N_{in}(T)=(1/2)\coth\left(\hbar\omega/2k_{B}T\right)$ is the input noise from the 50 $\mathrm{\Omega}$ terminations. $N_{qm}$ is the quantum noise and for $k_{B}T\ll\hbar\omega$ is equal to one half. $G_{PA},N_{A}$ and $G_{HEMT},N_{HEMT}$ are the gain and added noise of the KI-TWPA and the HEMT amplifier, respectively. $G_{w}$ is the additional effective warm amplification, and $N_{tot}$ is the total noise measured on the spectrum analyzer (SA).
When the KI-TWPA is pumped, the total noise measured by the SA can be calculated using equations (B1-B5).
@@
Then the added noise of the KI-TWPA is
@@
In the above equations, the loss factors $L_{1}$ and $L_{2}$ can be estimated either by measuring insertion losses of the components separately or using the bypass and pump-off transmission measurements using the relay switches. $N_{HEMT}$ can be estimated using either equation (B 12) or (B 11) with a y-factor measurement. Finally, measuring the gain of the KI-TWPA, we can calculate the added noise of the TWPA from the above equation. Using the unpumped KI-TWPA noise spectra, we measure an average HEMT noise temperature of 5 kelvin, which is in agreement with LNF-LNC0.3_14B datasheet.
In Fig. 5.b, you can see the transmission ($S_{21}$) measurements of the cable in two scenarios. The first one is the bypassing of KI-TWPA, represented by the orange line, and the second one is the transmission through DPX2, KI-TWPA, and DPX3, represented by the blue line. We have calculated the loss between switch B1 and the KI-TWPA by subtracting the unpumped measurement from the bypass measurement and dividing it by two since we used identical diplexers with similar insertion loss. This calculated loss is plotted as a green line in Fig. 5.b as a function of frequency, and we have used it as L1 and L2 loss factors in the main text.

### v4

@@ IV Conclusion
In conclusion, we have designed, fabricated, and measured a wide-bandwidth Four-Wave Mixing KI-TWPA that produces gain between 3 - 9 GHz with a 1 dB compression of -58 dBm suitable for the readout of large detector arrays and superconducting qubits. Using a Y-factor method, we have demonstrated a near quantum-limited noise performance of KI-TWPA in the 4.6 to 8 GHz frequency range. The amplifier produces gain in disjointed frequency ranges that we call the signal band and a higher-frequency idler (or image) band. This makes it straightforward to separate the idler tones using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies for the readout of cryogenic detectors. The signal band peak gain of this device is shown to be tunable with the frequency of the pump tone. This tunability makes the KI-TWPA suitable for applications where the signal frequency range needs to be adjusted, such as in dark matter experiments. In addition, the higher frequency idler band of the KI-TWPA can be used to look for dark matter candidates with a higher mass range. The simplest way to accomplish dual-band operation without changing the isolator and second-stage HEMT amplifier would be to operate the KI-TWPA in a frequency translating mode, relying on the fact that the idler band contains a redundant copy of the signal band and visa versa. A dual-band circuit could be implemented by adding two switches, one before the HP port of DPX1 and the other before the LP port of DPX2 to switch between either signal inputs or terminations. This capability makes it possible to cover a wide range of masses with a single KI-TWPA.
@@ Acknowledgements.
Figure 5: a) Circuit diagram of the noise measurement setup. The input line is attenuated using three cryogenic attenuators and then connected to switch C1. Two cryogenic 50 $\mathrm{\Omega}$ loads were used as noise sources. The ‘Hot’ noise source was placed at the 3 K stage of the dilution fridge, and the ‘Cold’ noise source was placed at the MXC. In this configuration, we can switch between the 50 $\mathrm{\Omega}$ for Y-factor measurements and the signal line for gain measurements. The Pump tone was injected in the KI-TWPA through a bandpass filter and DPX 1 and 2 with cross-over frequencies of 14 and 9 GHz, respectively. The pump tone is then routed out using DPX 3 (cross-over frequency of 9 GHz) and terminated on the 50 mK stage of the cryostat to prevent the HEMT from saturating. The cables used in the measurement setup connecting switch C1 to the ‘hot’ and ‘cold’ terminations on the 3 K and the MXC stages were NbTi superconducting cables. To calibrate the loss in the components, we used a cryogenic cable between switches B1 and B2 to bypass the TWPA and the diplexers. The output is then read out through an isolator, a cryogenic HEMT amplifier, and a room-temperature amplifier using a spectrum analyzer. b) The plot of measured transmission through the entire noise setup with the KI-TWPA unpumped is denoted as ‘Pump off,’ and through the cryogenic cable bypassing the components is denoted as ‘Bypass.’ The plot labeled ‘Loss’ is the difference between the two measurements divided by two to account for the loss in one set of diplexer and cable since identical components were used on both sides of the KI-TWPA.
@@ Appendix A Noise Measurement Setup
The pump input was filtered using a bandpass filter with a center frequency of 11.45 GHz and a bandwidth of 300 MHz at the MXC stage and directed to the KI-TWPA using DPX1 (Marki, DPX1114) and DPX2 (Minicircuits, ZDSS-7G10G-S+) as illustrated in Fig 5a. Using the highpass port of DPX3 (Minicircuits, ZDSS-7G10G-S+), the pump tone and the idler tones are terminated at the 50 mK stage of the dilution fridge. The total transmission of the system through the device was measured using a VNA with the pump tone off and the transmission of the cable bypassing the KI-TWPA and the diplexers as shown in Fig. 5.a. The plots comparing the transmission between the cryogenic cable and the unpumped KI-TWPA are shown in Fig. 5.b and labeled as ‘Bypass’ and ‘Pump off,’ respectively. The net loss from DPX2 and cables placed between the KI-TWPA and switch C1 in the circuit is also plotted in Fig. 5.b as ‘Loss.’
@@ Appendix B Noise Theory
To model the noise in our setup, we used the circuit diagram shown in Fig. 4.b. If we assume the mK losses in the components between the noise sources (hot & cold in this case) and the KI-TWPA are $L_{1}$ and $L_{2}$, the cascaded noise equations can be written as follows:
@@
$N_{tot}$ $=G_{w}\Big{(}N_{4}+N_{w}\Big{)}\ .$ (11)
Where $N_{in}(T)=(1/2)\coth\left(\hbar\omega/2k_{B}T\right)$ is the input noise from the 50 $\mathrm{\Omega}$ terminations. $N_{qm}$ is the quantum noise and for $k_{B}T\ll\hbar\omega$ is equal to one half. $G_{PA},N_{A}$ and $G_{HEMT},N_{HEMT}$ are the gain and added noise of the KI-TWPA and the HEMT amplifier, respectively. $G_{w}$ and $N_{w}$ are the additional effective warm amplification and warm amplifier noise, respectively. $N_{tot}$ is the total noise measured at the spectrum analyzer (SA).
When the KI-TWPA is pumped, the total noise measured by the SA can be calculated using equations (B1-B5). The overall noise contribution of the warm amplifier was ignored because of the high gain ($\sim$ 38 dB) of the HEMT amplifier and the KI-TWPA ($\sim$ 20 dB).
@@
Then, the added noise of the KI-TWPA is
@@
In the above equations, the loss factors $L_{1}$ and $L_{2}$ can be estimated either by measuring insertion losses of the components separately or using the bypass and pump-off transmission measurements using the relay switches. $N_{HEMT}$ can be estimated using either equation (B12) or (B11) with a y-factor measurement. Finally, measuring the gain of the KI-TWPA, we can calculate the added noise of the TWPA from the above equation. Using the unpumped KI-TWPA noise spectra, we measure an average HEMT noise temperature of 5 kelvin, which agrees with LNF-LNC0.3_14B datasheet.
In Fig. 5.b, we present the transmission ($S_{21}$) measurements of the cable in two scenarios. The first one is the bypassing of KI-TWPA, represented by the purple curve, and the second one is the transmission through DPX2, KI-TWPA, and DPX3, represented by the dark blue curve. We have calculated the loss between switch B1 and the KI-TWPA by subtracting the unpumped measurement from the bypass measurement and dividing it by two since we used identical diplexers with similar insertion loss. This calculated loss is plotted as an orange curve in Fig. 5.b as a function of frequency, and we have used it as L1 and L2 loss factors in the main text.